3. Consider rearranging the letters in the word FATHER (a) Find the number of 6 letter words that can be formed by considering all possible permutations of the letters in the word FATHER (b) How many of these words begin with F and end with R? (c ) What is the probability of forming a six letter word that begins with F and ends with R by randomly rearranging the letters in FATHER?
Answer #1

a) number of total permutation of 6 different alphabets =6! =720

b) number of such words =arrange remaining 4 alphabets =4! =24

c) probability =24/720 =1/30